OpenStudy (anonymous):

Write a quadratic equation with roots of 4 + 3i and 4 – 3i. Show all your work.

OpenStudy (welshfella):

this will be (x + (4 + 3i))( x + (4 - 3i)) = 0 expand this and you'l get your equation

OpenStudy (welshfella):

oops i made a mistake sorry it should be negatives after the x 's

OpenStudy (welshfella):

(x - (4 + 3i))( x - (4 - 3i)) = 0

OpenStudy (welshfella):

I'll start you off;- = x^2 - x(4 - 3i) -x(4 + 3i) +(4 + 3i)(4 - 3i) = 0

OpenStudy (welshfella):

recall that i^1 = -1

OpenStudy (welshfella):

x^2 - 4x + 3i - 4x - 3i + (16 - 12i + 12i- 3 i^2)

OpenStudy (welshfella):

now you can simplify this Note that the 3i' s and 12i's cancel out.